After a couple of years in Canada, you decide you want to start contributing to your retirement. How much money do you need to contribute at the end of each year for 37 years to accumulate to $213,000 in the RRSP account if interest is 8.04% compounded quarterly?
1
Expert's answer
2020-04-21T13:26:32-0400

"FV=PV\times\frac{(1+i)^n-1}{i}"


according to the postnumerando scheme

"213 000=PV\times\frac{(1+0.0201)^{148}-1}{0.0201}"


"213 000=PV\times\frac{18.02}{0.0201}"


PV=237.64


"237.64*4=950.56"